The capital accounts of the Low, Medium and High partnership at December 31, 2019, together with profit and loss sharing ratios are as follows:

                        Low (40%)                  204,000

                        Medium (20%)           118,000

                        High (40%)                278,000

The partners agree to admit Very High into the partnership.

REQUIRED: Prepare the journal entry or entries to admit Very High into the partnership under each of the following independent assumptions.

  1. Low sells half of his interest and medium sells 40% of her interest to Very High for a total purchase price of $190,000 paid directly to Low and Medium.
  2. Very High invests $180,000 cash in the partnership for a 20% interest in capital and profits and the partnership assets are not revalued.
  3. Very High invests $120,000 cash in the partnership for a 20% interest in capital and profits and the partnership assets are revalued.
  4. Very High invests $200,000 cash in the partnership for a 30% interest in the capital and profits and partnership assets are revalued.
Answer #1

Low sells half of his interest and medium sells 40% of her interest to Very High for a total purchase price of $190,000 paid directly to Low and Medium - Here no journal entries are required in the books of the partnership firm.   

Very High invests $180,000 cash in the partnership for a 20% interest in capital and profits and the partnership assets are not revalued

Existing partner capital =  600,000  (204,000+ 118,000+ 278,000)
New partner investment   =  180,000

Required Capital = 900,000 (180,000/20%)

The difference is treated as Goodwill

Goodwill = [900,000-(600,000+180,000)] = 120,000

This goodwill can be allocated to existing partners in the ratio 40:40:20

Therefore Journal entry is :

Account    Debit Credit
Cash 180,000
Goodwill 120,000
Capital – Partner Low 48,000
Capital – Partner Medium 24,000
Capital – Partner High 48,000
Capital – Partner Very High 180,000

Very High invests $120,000 cash in the partnership for a 20% interest in capital and profits and the partnership assets are revalued

Existing partner capital =  600,000  (204,000+ 118,000+ 278,000)
New partner investment   =  120,000

Required Capital = 600,000 (120,000/20%)

Therefore the Shortfall = [600,000 - (600,000+120,000)] = -120,000

720,000 x 20% = 144,000

144,000 - 120,000 = 24,000

It should be adjusted with the existing partners capital account in the ratio 40:20:40

Account    Debit Credit
Cash 120,000
Capital – Partner Low 9,600
Capital – Partner Medium 4,800
Capital – Partner High 9,600
Capital – Partner Very High 144,000

  

Very High invests $200,000 cash in the partnership for a 30% interest in the capital and profits and partnership assets are revalued

Existing partner capital =  600,000  (204,000+ 118,000+ 278,000)
New partner investment   =  200,000

Required Capital = 800,000

Very High's Capital Shoild be = 800,000 x 30% = 240,000

Difference in Capital of Very High = 240,000- 200,000 = 40,000

This can be adjusted to existing partners in the ratio 40:40:20

Therefore Journal entry is

Account    Debit Credit
Cash 200,000
Capital – Partner Low 16,000
Capital – Partner Medium 8000
Capital – Partner High 16,000
Capital – Partner Very High 240,000